If y(1) = – 3, y(3) = 9, y(4) = 30, and y(6) = 132, find the four-point

Lagrange interpolation polynomial that takes the same values as the function y at the given points.


1
Expert's answer
2021-01-06T19:53:01-0500
"\begin{matrix}
 x_0=1, & x_1=3, & x_2=4, & x_3=6 \\
 y_0=-3, & y_1=9, & y_2=30, & y_3=132
\end{matrix}"


The interpolating polynomial is:

"L(x)=\dfrac{(x-x_1)(x-x_2)(x-x_3)}{(x_0-x_1)(x_0-x_2)(x_0-x_3)}\times y_0"

"+\dfrac{(x-x_0)(x-x_2)(x-x_3)}{(x_1-x_0)(x_1-x_2)(x_1-x_3)}\times y_1"

"+\dfrac{(x-x_0)(x-x_1)(x-x_3)}{(x_2-x_0)(x_2-x_1)(x_2-x_3)}\times y_2"

"+\dfrac{(x-x_0)(x-x_1)(x-x_2)}{(x_3-x_0)(x_3-x_1)(x_3-x_2)}\times y_3"





"L(x)=\dfrac{(x-3)(x-4)(x-6)}{(1-3)(1-4)(1-6)}\times (-3)"

"+\dfrac{(x-1)(x-4)(x-6)}{(3-1)(3-4)(3-6)}\times 9"

"+\dfrac{(x-1)(x-3)(x-6)}{(4-1)(4-3)(4-6)}\times30"

"+\dfrac{(x-1)(x-3)(x-4)}{(6-1)(6-3)(6-4)}\times 132"

"=x^3-3x^2+5x-6"

"L(x)=x^3-3x^2+5x-6"